OpenStudy (anonymous):

original question: Simplify: (9x)^-2...I am confused...is the correct answer 1/9x^2 or is the correct answer 1/81x^2

OpenStudy (anonymous):

As mentioned earlier: (9^-2) = 1 / 81 (x^-2) = 1 / x2 If you'd multiply them together: (1/81)(1/x2) = (1/81x2)

OpenStudy (anonymous):

ok, ty

OpenStudy (anonymous):

express the answer in parenthesis as u have written?